# Mask as Supervision: Leveraging Unified Mask Information for Unsupervised 3D Pose Estimation

### v1

@@ 1 Introduction
Figure 2: Different acquisition methods for structural priors. Known as an important prior, structural information leads to plausible skeletons in pose estimation. Without considering it, most previous methods necessitate supervised post-processing (SPP) during inference. To tackle this issue, we propose a method that enables effective and ready-to-use pose estimation in a fully unsupervised fashion.
@@
Comprehensive experiments on Human3.6M [17] and MPI-INF-3DHP [33] datasets demonstrate that our method achieves superior performance among state-of-the-art unsupervised pose estimation methods. Additionally, our model possesses the capacity to leverage in-the-wild video data. Fine-tuning on TikTok dataset [18] without any kind of annotations, our model performs stronger generalization ability on unseen indoor and outdoor datasets.
@@ 2 Related Work
Unsupervised 2D Landmark Detection.
@@
Unsupervised 3D Pose Estimation.
@@
The closest setting to our method is [25], it adopts paired images with the same background and extracts texture and shape features from raw images and keypoints. It reaches unsupervised 3D pose estimation by applying a kinematic restriction in 3D and human prior from SMPL [30]. However, their results suffer from left-right ambiguity due to the lack of view constraints on the 2D mask representations.
As demonstrated later, our proposed method is fully unsupervised, requiring no annotations or templates, and effectively reduces left-right ambiguity.
@@ 3 Method
where $J$ represents the number of joints as defined by the particular dataset in use. Typically, the $\phi$ network is implemented via a backbone network [35], complemented by an underlying 3D heatmap representation. This heatmap subsequently predicts the 3D coordinates of the joints through an integral operation [49].
@@ 3.2 Mask as Supervision
Skeleton Mask from Body Structure Prior.
@@
As demonstrated in the experiments, the structure prior effectively addresses the majority of ambiguity issues within mask supervision. However, a gap still remains between the Skeleton Mask and the ground truth mask, resulting in the loss of fine details of the human body shape. The integration of a shape prior could further narrow this divergence, culminating in a more precise 3D pose estimation.
Physique Mask from Body Shape Prior.
@@
Geodesic Weighting for Hard Positives and Negatives.
It is well-established that joints further from the root on the kinematic tree are more challenging to estimate due to the increased variation [49]. Typically, these end joint positions appear in the mask’s locations that are distant from the centroid. Meanwhile, false positives lying on the background pixels should be unequally penalized according to the extent of deviation for smooth optimization. These insights motivate us to enhance the relevant hard positive and negative pixels on the mask ground truth $\mathbf{M^{gt}}$.
@@ 3.3 Leveraging Priors in Diverse Data Modalities
Video Data Modality.
Video is the most common visual data modality apart from images. It is comprised of a series of consecutive frames, where adjacent frames imply spatial and temporal consistency. Therefore, basic background modeling algorithms [3, 38, 7] and the emerging unsupervised methods based on motion [5, 27] allow us to efficiently extract the foreground mask from video data.
Multi-view Data Modality.
@@
Discussion: Ambiguity in Left-Right Reversal.
Owing to the symmetry of human body, using Masks as Supervision leads to a left-right reversal problem, also mentioned in [41, 11]. Networks may collapse that fail to distinguish the left and right side (i.e. all keypoints on the left of the image being treated as left joints) since both situations return the same error in the backward procedure.
In previous works, such as our Baseline in Eq. 3, the problem is neglected when introducing supervised post-processing (SPP) for evaluation. It utilizes annotations to train a linear or two hidden-layer network $\theta:\mathbb{R}^{L\times 3}\rightarrow\mathbb{R}^{J\times 3}$ mapping high dimensional keypoints (usually $L\geq 30$) to joints. Meanwhile, SPP makes it convenient to omit the processing of joint connectivity in their model design (further discussed in Sec. 4.6) while it leverages training set annotations thus not fully unsupervised.
In our framework, we adopt structure and shape prior to organizing human skeletons. To tackle the left-right reversal problem, we utilize the geometry constraints in Eq. 10 when multi-view data modality is accessible. The Skeleton Mask $\mathbf{M}_{Skel}^{\iota}$ is sketched by the re-projected 2D keypoints $\mathbf{x}_{prj}^{\iota}=\mathbf{\Pi}^{\iota}(\mathbf{X})$. Only by eliminating ambiguity in prediction, networks can receive correct keypoints from triangulation and projection procedures.
Table 1: Comparison with state-of-the-art methods on Human3.6M. SPP: supervised post-processing. UP: unpaired ground truth pose or its prior, T: manually designed template. SF: supervised flip to eliminate left-right ambiguity. indicates our results do not consider the ambiguity in left-right reversal. †† indicates we do not consider inner skeleton relationships and follow the common SPP settings. The best results in SPP and No SPP groups are marked in red and blue. MPJPEs are in $mm$.
@@
UP T SF Joints 2D MSE-MPJPE N-MPJPE P-MPJPE
@@
SPP ${}^{{\dagger}{\dagger}}$$\times$ $\times$ $\times$ 3D $2.52$ $\mathbf{65.5}$ $\mathbf{66.1}$ $\mathbf{61.9}$
@@
No SPP ${}^{{\dagger}}$$\times$ $\times$ 3D $3.17$ $\mathbf{85.6}$ $\mathbf{85.6}$ $\mathbf{79.3}$
@@ 4.2 Implementation Details
Mask Extraction.
General segmentation methods are widely used to extract foreground masks from video data. Human3.6M dataset provides the foreground mask by the off-the-shelf graph cut algorithm [3]. MPI-INF-3DHP dataset adopts chroma key to extract masks in the motion capture environment with a green screen.  [1] is utilized in TikTok dataset. With the help of the multi-prompt segmentation model, we select SAM [22] and obtain the mask with smooth and relatively accurate edges in our experiments.
Training Details.
Experimentally, Batch Normalization [16] provides inaccurate batch statistics estimation during inference, leading to poor generalization capabilities in resolving ambiguities related to left-right reversal. Hence, we adopt Group Normalization [55] to replace Batch Normalization in our detector, which avoids parameter estimation during inference.
@@ 4.3 Quantitative Analysis
Figure 5: Qualitative results on Human3.6M and MPI-INF-3DHP datasets. We visualize joints in both 2D and 3D coordinate systems, for Human3.6M (left) and MPI-INF-3DHP (right) datasets. Every $4^{th}$ column in each sub-figure shows 3D ground truth joints.
Results on Human3.6M.
In Tab. 1, we compare against both 2D and 3D unsupervised pose estimation methods on the Human3.6M dataset. In previous studies, several laborious human-involved annotations or priors are introduced, which deviate from the strict definition of unsupervised learning. To ensure a fair comparison and clarify the experimental settings, we succinctly summarized them as follows: SPP (supervised regression in post-processing), SF (supervised flip to eliminate left-right ambiguity), UP (unpaired pose or its prior), T (manually designed template).
@@
Results on MPI-INF-3DHP.
With the aim to evaluate the performance on challenge scenarios, we conduct our model on MPI-INF-3DHP dataset [33], which involves a larger diversity of actions for further exploration. Tab. 2 presents our contestable performance with unpaired labels or SPP-necessitated methods. Detailly, without any kind of annotations, we achieve advanced performance in MPJPE and considerable accuracy in PCK and AUC metrics. Note that not many unsupervised methods directly experiment on this challenging dataset. We demonstrate our results under the suggested PCK and AUC metrics in [33] calling for follow-up experiments.
@@ 4.5 Leveraging In-the-wild Data
Vast amounts of data from various sources are being generated, including real-world scenarios such as video sequences captured by a single camera. As a result, it is crucial and essential for unsupervised models that can effectively learn from in-the-wild data to enhance model robustness and generalization ability. Under our circumstances, our model possesses the ability to take single-view images as input, where the foreground masks can be extracted in the same way as multi-view video data.
Figure 6: Qualitative results on TikTok dataset. We visualize the Mixed-view fine-tuning predictions in different views.
Table 3: Evaluations with different training strategies. MPJPEs are in $mm$. Note models are not trained on 3DHP [33] and MPII [2] datasets.
Settings H3.6M 3DHP MPII
MPJPE($\downarrow$) PCK($\uparrow$) PCKh($\uparrow$)
Baseline $101.63$ $60.1$ $26.7$
Multi-view ft $95.90$ $61.2$ $27.3$
Mixed-view ft $94.82$ $63.4$ $28.1$
improvement +1.08 +2.2 +0.8
To evaluate our model capacity of learning from in-the-wild data and to overcome the shortages of appearance diversity in mocap datasets, we conduct on TikTok dataset [18], which provides single-view videos without any pose annotations for view mix training. Formally, a fully initialized model on Human3.6M is chosen as the baseline. Then we train it on both Human3.6M and TikTok datasets, named Mixed-view Fine-tuning. For comparison, we take another strategy that merely uses Human3.6M data to train for the same epochs, named Multi-view Fine-tuning. We evaluate them on seen and unseen datasets, containing indoor and outdoor scenarios. The qualitative and quantitative results are demonstrated in Fig. 6 and Tab. 3.
It is insightful that Mix-view Fine-tuning surpasses Multi-view Fine-tuning in every metric, indicating the capability of absorbing in-the-wild data in our framework.
@@ 4.6 Ablation Study
Skeleton Mask as Structural Representation.
@@
Physique Mask as Shape Representation.
@@
Although without triangulation, the model gains better performance in MPJPE, it raises large ambiguity in left-right reversal according to our metric $r$.
@@ 5 Conclusion
We present an unsupervised 3D pose estimation framework that leverages unified mask information from structure and shape representations. The proposed model effectively organizes the human skeleton with priors from the mask, making it capable of providing interpretable keypoints and accessing annotation-free data. Performance on two widely used datasets demonstrates the superior ability of our model. Moreover, we illustrate the model can be enhanced by in-the-wild video data, by virtue of the fully unsupervised manner.
@@ Acknowledgement
Supplementary Material
Appendix A Cascade Training
A.1 Auxiliary Loss
Apart from mask reconstruction loss in the Sec. 3.2, several auxiliary losses are introduced to strengthen the unsupervised system.
Constant bone length loss.
Pre-define a list of bone lengths, including torso-pelvis, thorax-torso, neck-thorax, nose-neck, and head-nose, denoted as $C$. We utilize a sub-optimal loss to initialize a general skeleton.
$L_{bone_{c}}=\lambda_{bone_{c}}\sum_{b}^{B}||\mathbf{B}_{b}-C_{b}||_{2}^{2}$ (12)
Equivalance loss.
Same as [44], we expect keypoints with spatial invariance can be discovered. Let denote $\mathcal{T}$ as a 2D affine transformation, $\phi$ as our keypoint detector, and $\mathbf{I}$ as the input image. The model is constrained to predict keypoints with equivalence constraints that enforce the commutativity of applying a transformation $\mathcal{T}$ either before or after detection.
$L_{eq}=\lambda_{eq}\sum_{i=0}^{J}||\mathcal{T}(\phi(\mathbf{I}))-\phi(\mathcal{T}(\mathbf{I}))||_{2}^{2}$ (13)
Symmetry loss.
Considering the inner symmetry of the human skeleton, we leverage this prior to providing supervision. In detail, the shoulders and hips are bilaterally symmetrical to the thorax and pelvis, respectively. We denote $\mathbf{x}_{l}$ and $\mathbf{x}_{r}$ as the left and right side of endpoints, $\mathbf{x}_{c}$ as symmetrical midpoints in 2D.
$L_{sym}=\lambda_{sym}||(\mathbf{x}_{l}+\mathbf{x}_{r})/2-\mathbf{x}_{c}||_{2}^{2}$ (14)
A.2 Bone Mask.
As mentioned in the body structure prior, each bone mask $\mathbf{B}_{b}$ is formally defined as
where $\sigma$ is a hyper-parameter controlling bone width, and $d_{b}$ is the $L_{2}$ distance between pixel $\mathbf{p}$ in the map and the line segment. Specifically, the distance $d_{b}$ is defined as
$d_{b}=\left\{\begin{cases}||\mathbf{p}-\mathbf{x}(i_{b})||_{2},&\text{if $r\leq 0$}\\ ||\mathbf{p}-(\mathbf{x}(i_{b})+r\mathbf{x}(j_{b}))||_{2},&\text{if $0<r<1$,}\\ ||\mathbf{p}-\mathbf{x}(j_{b})||_{2},&\text{if $r\geq 1$,}\end{cases}\right.$ (16)
where,
$r=\frac{(\mathbf{p}-\mathbf{x}(i_{b}))\cdot(\mathbf{x}(i_{b})-\mathbf{x}(j_{b}))}{||\mathbf{x}(i_{b})-\mathbf{x}(j_{b})||_{2}^{2}}.$ (17)
A.3 Training Details
To gradually approach the ground truth and avoid numerous local optima in training, we design a cascade training procedure to disentangle bones. Different connections in the skeleton are demonstrated in Fig. 7.
(a) Tight.
(b) Medium.
(c) Loose.
Figure 7: Schematic of different skeleton designs. Green lines indicate auxiliary connections for better mask fitting.
For the first stage, we utilize the tight form to depict the skeleton mask $\mathbf{M}_{Skel}$. It selects dense connections in the torso and omits the connections at the arms since the main part shows less diversity which is beneficial for the model to warm up. In initialization, we first take the gravity centroid and a point with a fixed distance from the centroid as pseudo pairs to supervise the predicted pelvis and thorax for $400$ steps. Following that, two pivotal joints are coarsely located. Meanwhile, Constant bone length loss $L_{bone_{c}}$, Equivalance loss $L_{eq}$, and Symmetry loss $L_{sym}$ are adopted to assist the reconstruction of the skeleton mask $L_{Skel}$. In detail, we take $\lambda_{s}=10.0$ with geodesic weighting and $2.5,1.0,2.0$ for $\lambda_{bone_{c}},\lambda_{eq},\lambda_{sym}$ respectively. Note that $\lambda_{bone_{c}}$ is a sub-optimal target, the given $\lambda_{bone_{c}}$ will only generate a small loss in the order of $31e-$ 3. With geodesic weighting, the selected joints are progressively driven to ideal locations from the gathered. This procedure takes $15$ epochs with a $42e-$ 4 learning rate.
After the reconstruction loss in the skeleton mask meets convergence, we then adopt the physique mask $\mathbf{M_{Physo}}$ to replenish mask representations. In this stage, we aim to train the Shape Prior Network $\psi$ to drive the arm joints in the final stage. The medium form of the skeleton mask is leveraged which contains fewer auxiliary connections and provides more flexibility in pose estimation. In terms of loss function, we remove the Constant bone length loss $L_{bone_{c}}$ while keeping the Equivalance loss $L_{eq}$ and Symmetry loss $L_{sym}$. To avoid the unexpected extension of shoulders (i.e. extend to the arms region), we remove the geodesic weighting for both the skeleton mask and physique mask and take $1.0$ and $4.0$ for $\lambda_{s}$ and $\lambda_{p}$. Besides, we force both the left and the right hand, as well as the left and the right elbow, to be positioned in the same location. It eliminates bias between the left and right parts as an initialization for the last stage. Experimentally, we constrained it for the first $600$ steps. The Shape Prior Network undertakes a simple task and converges fast in $5$ epochs.
In the last stage, we fully utilize the proposed modules, including the skeleton mask, the physique mask, and geodesic weighting. The loose form opts to offer enough flexibility to estimate the locations of the arm parts. We keep the auxiliary loss in the second stage and take $2.0$ and $1.5$ for $\lambda_{s}$ and $\lambda_{p}$. The entire framework is trained with a multi-step schedule that the learning rate starts at $42e-$ 4 for the first $14$ epochs and decays to $41e-$ 4 for the next $11$ epochs.
The bone width $\sigma$ in the skeleton mask is set round $3\times 10^{-3}$ depending on different cascade periods. We use Adam optimizer [21] with a learning rate $2\times 10^{-4}$. The network is trained 45 epochs in total with batch size 32. It takes about 9 hours on two 3090 GPUs.
Appendix B Ablation Study
In this section, we supplemented qualitative and quantitative results in the ablations study to further elaborate on the effectiveness of our proposed components.
B.1 Skeleton mask as structural representation.
Note that our structure prior can degenerate into a Gaussian-like point-based representation. Therefore, with configuration hyperparams controlled, we practically implement the baseline method by taking joints and themselves as bone connections.
Table 5: Ablation study on the skeleton mask. $\psi$: shape prior network. SPP: supervised post-processing. We remove skeleton priors to demonstrate the consequences of the lack of structural information.
Configurations 2D MSE-($\downarrow$)
wo structure prior, wo $\psi$ $20.22$
wo structure prior $20.06$
wo structure prior + SPP $9.32$
structure prior & $\psi$ $3.63$
The quantitative results in Tab. 5 demonstrate the effectiveness of structure prior and the utilities to alternate the supervised post-processing (SPP), which is compatible with the conclusion in Sec. 4.6. Formally, compared with the first two configurations, it indicates the structural prior is learned from our skeleton mask and the physique Mask will not provide structural information since the generator $\psi$ treats all keypoints as equal. After adding the powerful but supervised SPP, the performance grows with the post-processing network learned from training poses. In the last line of configurations, with structure prior, the plausible skeletons can be constructed without supervision and present lower error.
B.2 Physique Mask as Shape Representation.
Figure 8: Visualizations of the configurations that remove shape prior network and geodesic weighting.
In Fig. 8, we visualize the results without shape prior network and geodesic weighting. It can be seen that keypoints are located in the main part of the body since the upper body shows larger spatial diversity causing problems in optimization. The Shape Prior Network $\psi$ provides shape-level representations, requiring keypoints expansion as hints for a complete reconstructed mask and eventually drives keypoints to the extremities. Moreover, with the help of geodesic weighting, which constructs a smoother solution space and deals with corner cases, leading to further improved performance.

### v2

@@ Abstract
Keywords:
Unsupervised Learning Pose Estimation Mask
@@ 1 Introduction
Figure 2: Different acquisition methods for structural priors. Known as an important prior, structural information leads to plausible skeletons in pose estimation. Without considering it, most previous methods necessitate supervised post-processing (SPP) during inference. To tackle this issue, we propose a method that enables effective and ready-to-use pose estimation in an unsupervised fashion.
Comprehensive experiments on Human3.6M [17] and MPI-INF-3DHP [32] datasets demonstrate that our method achieves superior performance among state-of-the-art unsupervised pose estimation methods. Additionally, our model possesses the capacity to leverage in-the-wild video data without any kind of annotations. With increased training data, our model performs gradually strengthened generalization ability on unseen outdoor datasets.
@@ 2 Related Work
2.0.1 Unsupervised 2D Landmark Detection.
@@
2.0.3 Unsupervised 3D Pose Estimation.
@@
The closest setting to our method is [24], it adopts paired images with the same background and extracts texture and shape features from raw images and keypoints. It reaches unsupervised 3D pose estimation by applying a kinematic restriction in 3D and human prior from SMPL [29]. However, the unpaired 3D pose prior is still necessitated and remains a gap to free the supervision from pose annotation.
As demonstrated later, our proposed method requires no pose annotations or templates and effectively organizes skeletal information for interpretable joint prediction.
@@ 3 Method
where $J$ represents the number of joints as defined by the particular dataset in use. Typically, the $\phi$ network is implemented via Hourglass [34] as the backbone network, complemented by an underlying 3D heatmap representation. This heatmap subsequently predicts the 3D coordinates of the joints through an integral operation [48].
@@ 3.2 Mask as Supervision
3.2.3 Skeleton Mask from Body Structure Prior.
@@
As demonstrated in the experiments, the structure prior builds a connection from keypoint vectors to shape maps and effectively extracts skeletal information of the majority part within mask supervision. However, a gap still remains between the Skeleton Mask and the ground truth mask, resulting in the loss of fine details of the human body shape. The integration of a shape prior could further narrow this divergence, culminating in a more precise 3D pose estimation.
3.2.4 Physique Mask from Body Shape Prior.
@@
3.2.5 Geodesic Weighting for Hard Positives and Negatives.
It is well established that joints further from the root on the kinematic tree are more challenging to estimate due to the increased variation [48]. Typically, these end joint positions appear in the mask’s locations that are distant from the centroid. Meanwhile, false positives lying on the background pixels should be unequally penalized according to the extent of deviation for smooth optimization. These insights motivate us to enhance the relevant hard positive and negative pixels on the mask ground truth $\mathbf{M^{gt}}$.
@@ 3.3 Leveraging Priors in Diverse Data Modalities
3.3.1 Video Data Modality.
Video is the most common visual data modality apart from images. It is comprised of a series of consecutive frames, where adjacent frames imply spatial and temporal consistency. Therefore, basic background modeling algorithms [3, 37, 7] and the emerging motion-based unsupervised methods [5, 26] allow us to efficiently extract the foreground mask from video data.
3.3.2 Multi-view Data Modality.
@@
3.3.3 Discussion: SPP Usage and Ambiguity in Left-Right Reversal.
Owing to the symmetry of human body, using Masks as Supervision leads to a left-right reversal problem, also mentioned in [40, 12]. Networks may collapse that fail to distinguish whether a person is facing or back to the camera (i.e. all keypoints on the left of the image being treated as left joints) since both situations return the same error in the backward procedure.
In previous supervised post-processing (SPP) related works, such as our Baseline in Eq. 3, they can predict landmarks and some locate near to ground truth (e.g. the first line of Fig. 2). However, physically interpretable joints cannot be obtained since they treat joints equally and omit connectivity. Therefore, in evaluation, SPP utilizes annotations to train a linear or two hidden-layer network $\theta:\mathbb{R}^{L\times 3}\rightarrow\mathbb{R}^{J\times 3}$ mapping high dimensional landmarks (usually $L\geq 2\times J$) to joints. The reversal problem is neglected with such SPP mapping. Worse still, it leverages training set annotations thus not fully unsupervised.
In our framework, we adopt structure and shape priors to organizing human skeletons and do not build on annotations in both training and evaluation. To tackle the left-right reversal problem, we utilize the geometry constraints in Eq. 10 when multi-view data modality is accessible. The Skeleton Mask $\mathbf{M}_{Skel}^{\iota}$ is sketched by the re-projected 2D keypoints $\mathbf{x}_{prj}^{\iota}=\mathbf{\Pi}^{\iota}(\mathbf{X})$. It confines that different views correspond to the same 3D pose. Only by eliminating ambiguity in each view, networks can obtain correct 3D keypoints from triangulation.
3.4 Cascaded Optimization
@@
4 Experiments
[…]
@@ 4 Experiments
UP T SF Joint MSE MPJPE N-MPJPE P-MPJPE
@@
SPP Ours†† $\times$ $\times$ $\times$ 3D $2.52$ $\mathbf{65.5}$ $\mathbf{66.1}$ $\mathbf{61.9}$
@@
No SPP Ours $\times$ $\times$ 3D $3.17$ $\mathbf{85.6}$ $\mathbf{85.6}$ $\mathbf{79.3}$
@@ 4.2 Implementation Details
4.2.1 Mask Extraction.
General segmentation methods are widely used to extract foreground masks from video data. Human3.6M dataset provides the foreground mask by the off-the-shelf graph cut algorithm [3]. MPI-INF-3DHP dataset adopts chroma key to extract masks in the motion capture environment with a green screen.  [1] is utilized in TikTok dataset. With the help of the multi-prompt segmentation model, we select SAM [21] to fine-tune the mask with smooth and relatively accurate edges in our experiments.
4.2.2 Training Details.
Experimentally, Batch Normalization [16] provides inaccurate batch statistics estimation during inference, leading to poor generalization capabilities in resolving ambiguities related to left-right reversal. Hence, we adopt Group Normalization [54] to replace Batch Normalization in our detector, which avoids parameter estimation during inference. More details can be found in the appendix.
@@ 4.3 Quantitative Analysis
4.3.1 Results on Human3.6M.
In Tab. 1, we compare against both 2D and 3D unsupervised pose estimation methods on the Human3.6M dataset. In previous studies, several laborious human-involved annotations or priors are introduced, which deviate from the strict definition of unsupervised learning. To clarify the experimental settings, we succinctly summarized them as follows: SPP (supervised regression in post-processing), SF (supervised flip to eliminate left-right ambiguity), UP (unpaired pose or its prior), T (manually designed template).
@@
Figure 5: Qualitative results on Human3.6M and MPI-INF-3DHP datasets. We visualize joints in 2D and 3D coordinate systems, for Human3.6M (left) and MPI-INF-3DHP (right) datasets. Every $4^{th}$ column in each sub-figure shows 3D ground truth joints.
4.3.2 Results on MPI-INF-3DHP.
To evaluate the performance on challenging scenarios, we conduct our model on MPI-INF-3DHP dataset [32], which involves a larger diversity of actions for further exploration. Tab. 2 presents our contestable performance with unpaired labels or SPP-necessitated methods. Detailly, without any kind of annotations, we achieve advanced performance in MPJPE and considerable accuracy in PCK and AUC metrics. Note that not many unsupervised methods directly experiment on this challenging dataset. We demonstrate our results under the suggested PCK and AUC metrics in [32] calling for follow-up experiments.
@@ 4.5 Leveraging In-the-wild Data
Vast amounts of data from various sources are being generated, including real-world scenarios such as video sequences captured by a single camera. As a result, it is crucial and essential for unsupervised models that can effectively learn from in-the-wild data to enhance model robustness and generalization ability. With the ability to process in-the-wild data, we firstly conduct a comprehensive analysis of the relationship between data amount and model performance in unsupervised pose estimation.
Scalability. To evaluate the scalability of the proposed mask as supervision, we collect all hands-on data with mask, including Human3.6M, MPI-INF-3DHP and in-the-wild TikTok [18] without any pose annotations. In Fig. 6, the model is trained step-by-step with scaled-up data and evaluated on the in-the-wild MPII [2]. With clear improvements, it demonstrates the capacity and scalability of our proposed method.
Figure 6: Scalability and generalization ability analysis. The detector is trained in mask and full supervision with training data scaling up. Sub-figures show the evaluation results on MPII and Human3.6M.
Figure 7: Qualitative results on TikTok dataset. We visualize the predictions from different views.
Generalization. Additionally, the supervised can only be trained with 3D annotation when multi-view data is accessible. Keeping the detector unchanged, the fully supervised detector on in-door Human3.6M cannot generalize as effectively to in-the-wild MPII as mask supervised one (from the right sub-figure to the left). Qualitatively, the model that leverages in-the-wild data with mask as supervision can overcome the shortages of appearance diversity in MoCap datasets. The visualization of TikTok test split is in Fig. 7.
@@ 4.6 Ablation Study
4.6.1 Skeleton Mask as Structural Representation.
@@
4.6.2 Physique Mask as Shape Representation.
@@
Although without triangulation $\Delta$, the model gains better performance in MPJPE, it raises large ambiguity in left-right reversal according to our metric $r$. In summary, each component of the physique mask plays a positive role in the whole framework and can be united effectively.
@@ 5 Conclusion
We present an unsupervised 3D pose estimation framework that leverages unified mask information from structure and shape representations. The proposed model effectively organizes the human skeleton with priors from the mask, making it capable of providing interpretable keypoints and accessing annotation-free data. Performance on two widely used datasets demonstrates the superior ability of our model. Moreover, we illustrate the model can be enhanced by in-the-wild video data, by virtue of the mask as supervision manner.